# Parameter Symmetry Breaking and Restoration Determines the Hierarchical Learning in AI Systems

### v1

@@ A.4 Neural Collapse
Figure 15: The raw data for Figure 6, but two networks have different number of neurons. The legend $(i,j)$ denotes the alignment between the $i$-th layer of network A and the $j$-th layer of network B. Left: linear network. Right: tanh network.
In Figure 4, we train ResNet18 on the CIFAR-10 dataset using vanilla weight decay and syre (Ziyin, 2024). In both cases, the weight decay is set to $5\times 10^{-4}$. Networks are trained for $200$ epochs, and results from the final epoch are reported. To generate Figure 4, we randomly select 10 images per class and compute the correlation between their features (i.e., the input to the final layer). The corresponding eigenvalue spectrum is presented in Figure 11, showing 10 prominent eigenvalues. Figures 12 and 13 extend these experiments with a weight decay of $5\times 10^{-3}$, where the vanilla weight decay model exhibits a stronger low-rank structure compared to the syre model.
@@
In Figure 6, we train two deep linear networks on MNIST using the Adam optimizer. Each network is a six-layer fully connected network (FCN) with 128 neurons per hidden layer. The average alignment to input data is computed as the average alignment between the data and the third layer during training. Figure 15 explores a similar setting for both linear and tanh networks, and the two networks have hidden layers with $64$ and $128$ neurons, respectively, where the universal alignment effects still hold.
@@ B.1 Formal Statement of Theorem 1
The second part of the theorem means that in the kernel regime [^2] Technically, this is the lazy training limit (Chizat et al., 2018)., being at a symmetric solution implies that the feature kernel features are being masked by the projection matrix $\nabla_{\theta}f(x,\theta_{0})\to P_{G}\nabla_{\theta}f(x,\theta_{0})$, and learning can only happen given these masks. The proof is a slight generalization of Propostions 2 and 3 in (Ziyin et al., 2025b).
@@
B.2 Formal Statement of Theorem 2
We will consider training a deep linear network with SGD and MSE loss
$\ell(\theta,x)=||W_{D}\cdots W_{1}x-y(x)||^{2},$ (26)
on datasets $\mathcal{D}_{M}=\{(Mx_{i},y_{i})\}_{i}$, where $M$ is an invertible matrix, and $y_{i}=Vx_{i}+\epsilon_{i}$ for i.i.d. noise $\epsilon_{i}$. We make the same assumptions as Theorem 5.4 of (Ziyin et al., 2024) (the most important of which is the SDE approximation to the SGD). Also, we use the definition of the noise equilibrium in (Ziyin, 2024), which essentially means that SGD reaches stationarity in the degenerate directions of the double rotation symmetry.
@@ B.2 Formal Statement of Theorem 2
Let $V^{\prime}=\sqrt{\Sigma_{\epsilon}}V\sqrt{\Sigma_{x}}$, $\text{rank}(V^{\prime})=d$ and $S^{\prime}$ be a diagonal matrix containing singular values of $V^{\prime}$. Consider two deep linear networks A and B with weights of arbitrary dimensions larger than $d$. Let model A train on $\mathcal{D}_{M}$ and model B on $\mathcal{D}_{M^{\prime}}$. Then, at the global minimum and at the noise equilibrium, every hidden layer of A is perfectly aligned with every hidden layer of B for any $x$, in the sense that
$h_{A}^{L_{A}}(x)=c_{0}Rh_{B}^{L_{B}}(x)$ (27)
for $1\leq L_{A}<D_{A}$ and $1\leq L_{B}<D_{B}$ and any $x$, where $c_{0}=\left(\frac{{\rm Tr}S^{\prime}}{d}\right)^{\frac{2L_{A}-D_{A}}{2D_{A}}-\frac{2L_{B}-D_{B}}{2D_{B}}}$ is a constant and $R=U_{1}U_{2}^{T}$, satisfying $U_{1}^{T}U_{1}=U_{2}^{T}U_{2}=I_{d}$. $h_{A}^{L_{A}}(x):=\Pi_{i=1}^{L_{A}}W_{i}^{A}Mx$, $h_{A}^{L_{A}}(x):=\Pi_{i=1}^{L_{A}}W_{i}^{A}M^{\prime}x$ denote the output of the $L_{A},L_{B}-$the layer of network A and B, respectively.
Proof.
Let $V^{\prime}:=\tilde{U}S^{\prime}\tilde{V}$ be its SVD. According to Theorem 5.4 of (Ziyin et al., 2024). At the global minimum and noise equilibrium under SGD, the solution of a $D_{A}$-layer network for the dataset $\mathcal{D}_{M}$ is given by
$\sqrt{\Sigma_{\epsilon}}M_{2}W^{A}_{D_{A}}=\tilde{U}\Sigma_{D}U_{D_{A}-1}^{T},\ W^{A}_{i}=U_{i}\Sigma_{i}U_{i-1}^{T},\ W^{A}_{1}M_{1}\sqrt{\Sigma_{x}}=U_{1}\Sigma_{1}\tilde{V}$ (28)
for $i=2,\cdots,D-1$, where $U_{i}$ are arbitrary matrices satisfying $U_{i}^{T}U_{i}=I_{d\times d}$, and
$\Sigma_{1}=\Sigma_{D}=\left(\frac{d}{{\rm Tr}S^{\prime}}\right)^{(D_{A}-2)/2D_{A}}\sqrt{S^{\prime}},\ \Sigma_{i}=\left(\frac{{\rm Tr}S^{\prime}}{d}\right)^{1/D_{A}}I_{d\times d}.$ (29)
We can verify that $\Pi_{i=1}^{D_{A}}W_{i}^{A}M=V$, or $h_{A}^{D_{A}}(x)=Vx$ as expected. The solution suggests that
$\Pi_{i=1}^{L_{A}}W_{i}^{A}M=U_{L_{A}}\left(\frac{{\rm Tr}S^{\prime}}{d}\right)^{\frac{2L_{A}-D_{A}}{2D_{A}}}\sqrt{S^{\prime}}\tilde{V}.$ (30)
Similarly
$\Pi_{i=1}^{L_{B}}W_{i}^{B}M^{\prime}=U_{L_{B}}\left(\frac{{\rm Tr}S^{\prime}}{d}\right)^{\frac{2L_{B}-D_{B}}{2D_{B}}}\sqrt{S^{\prime}}\tilde{V}.$ (31)
The proof is complete by comparing (30) and (31). ∎
B.3 Space Quantization Conjecture
Theorem 5.
Consider the loss $\ell(\theta)=\ell_{0}(\theta)+\gamma||\theta||^{2}$ with $\theta:=(\theta_{1},\cdots,\theta_{k})$ and $\theta_{i}\in\mathbb{R}^{n}$ for $1\leq i\leq k$. Assume that $\ell_{0}(\theta_{1},\cdots,\theta_{k})$ has the permutation symmetry $\ell_{0}(\theta_{1},\cdots,\theta_{i},\cdots,\theta_{j},\cdots,\theta_{k})=\ell_{0}(\theta_{1},\cdots,\theta_{j},\cdots,\theta_{i},\cdots,\theta_{k})$ for any $1\leq i,j\leq k$ and satisfies the following $q-$Lipschitz condition
$||\nabla\ell_{0}(\theta)-\nabla\ell_{0}(\theta^{\prime})||\leq K||\theta-\theta^{\prime}||^{q}$ (32)
for $q\geq 1$. Moreover, assume that $\inf_{\theta}\ell_{0}(\theta)>-\infty$ and that $K$ scales with the number of active neurons $m$ as $K=K_{0}m^{-\alpha}$. Then, at any global minimum,

### v2

@@ A.4 Representation Learning
Figure 15: An extension of Figure 6, where two networks have different number of neurons. The legend $(i,j)$ denotes the alignment between the $i$-th layer of network A and the $j$-th layer of network B. Left: linear network. Right: tanh network.
In Figure 5, we train ResNet18 on the CIFAR-10 dataset using vanilla weight decay (loss $\ell(\theta)+\gamma||\theta||^{2}$) and syre (loss $\ell(\theta)+\gamma||\theta-\theta_{0}||^{2}$) [83]. In both cases, the weight decay $\gamma$ is set to $5\times 10^{-4}$. Networks are trained for $200$ epochs, and results from the final epoch are reported. To generate Figure 5, we randomly select 10 images per class and compute the correlation between their features (i.e., the input to the final layer). The corresponding eigenvalue spectrum is presented in Figure 11, showing 10 prominent eigenvalues. Figures 12 and 13 extend these experiments with a weight decay of $5\times 10^{-3}$, where the vanilla weight decay model exhibits a stronger low-rank structure compared to the syre model.
@@
In Figure 6, we train two deep linear networks on MNIST using the Adam optimizer, where the input of the second one is randomly transformed by an invertible matrix, , whose elements are sampled from the standard Gaussian distribution. Each network is a six-layer fully connected network (FCN) with 128 neurons per hidden layer. The average alignment to input data is computed as the average alignment between the data and the third layer during training. Figure 15 explores a similar setting for both linear and tanh networks, and the two networks have hidden layers with $64$ and $128$ neurons, respectively, where the universal alignment effects still hold.
@@ B.1 Formal Statement of Theorem 1
The second part of the theorem means that in the kernel regime [^1] Technically, this is the lazy training limit [13]., being at a symmetric solution implies that the feature kernel features are being masked by the projection matrix $\nabla_{\theta}f(x,\theta_{0})\to P_{G}\nabla_{\theta}f(x,\theta_{0})$, and learning can only happen given these masks. The proof is a slight generalization of Propostions 2 and 3 in Ref. [88].
@@
B.2 Space Quantization Conjecture
@@
Consider the loss $\ell(\theta)=\ell_{0}(\theta)+\gamma||\theta||^{2}$ with $\theta:=(\theta_{1},\cdots,\theta_{k})$ and $\theta_{i}\in\mathbb{R}^{n}$ for $1\leq i\leq k$. Assume that $\ell_{0}(\theta_{1},\cdots,\theta_{k})$ has the permutation symmetry $\ell_{0}(\theta_{1},\cdots,\theta_{i},\cdots,\theta_{j},\cdots,\theta_{k})=\ell_{0}(\theta_{1},\cdots,\theta_{j},\cdots,\theta_{i},\cdots,\theta_{k})$ for any $1\leq i,j\leq k$ and satisfies
$||\nabla_{\theta}\ell_{0}(\theta_{1},\cdots,\theta_{i}+\theta,\cdots,\theta_{j}-\theta,\cdots,\theta_{k})||\leq K||\theta||^{q}$ (26)
for $q\geq 1$ and any pair of neurons. Moreover, assume that $\inf_{\theta}\ell_{0}(\theta)>-\infty$ and that $K$ scales with the number of active neurons $m$ as $K=K_{0}m^{-\alpha}$. Then, at any global minimum,